write each equation in slope-intercept form:
x/3+y/4=1
2x-3y=12

Answer by stanbon(75887)

write each equation in slope-intercept form:
x/3+y/4=1
Multiply thru by 12:
4x + 3y = 12
----
solve for "y":
y = (-4/3)x + 4
------------------------
2x-3y=12
---
3y = 2x-12
y = (2/3)x - 4
=======================
Cheers,
Stan H.
=====================